OpenStudy (ny,ny):

Hey!!! I have a quiz 2moro =( can u help me study? im in seventh grade so im no big deal really, considering all u ppl r in high school lol. O.O i alredy got a problem! a/5 + 8 = 9! its not tht hard, but its hard fer me since i take 9th grade math, im skipping 8th grade algebra!! Cathcing up is hard, try to help, k?

OpenStudy (lilg132):

a/5 + 8 = 9 a/5 = 9 - 8 a/5 = 1 a/5 * 5 = 1 * 5 (we do this so we get a by it self to work out what a is) a = 5

OpenStudy (lilg132):

dont forget to click good answer :p

OpenStudy (ny,ny):

oh, so we're supposed to multiply it by the 5? didnt kno tht lol

OpenStudy (ny,ny):

thank you!

OpenStudy (ny,ny):

do u wanna stay and help me study?

OpenStudy (lilg132):

u multipy it by 5 so you get a by it self in order to solve what the value of a is your welcome

OpenStudy (lilg132):

just write out the questions again like you did with this and im sure me or someone else will help you

OpenStudy (ny,ny):

kk

OpenStudy (ny,ny):

haha i thought u multiply it by 1, but tht was kinda stupid..

OpenStudy (ny,ny):

=)

OpenStudy (agreene):

\[\frac{a}5 + 8 = 9!\] \[a= 1814360\] ;)

OpenStudy (ny,ny):

tht can not be correct!

OpenStudy (lilg132):

write out your next question seperately not on this same page

OpenStudy (ny,ny):

ok

OpenStudy (agreene):

a/5 + 8 = 9! a/5 = 9! - 8 a = 5( 9! - 8) a= 5( (9*8*7*6*5*4*3*2*1) -8 ) a = 5( 362880 - 8 ) a = 5 ( 362872) a = 1814360

OpenStudy (ny,ny):

if u say this, then u r saying 1814360 / 5 + 8 = 9, tht is not true

OpenStudy (ny,ny):

the answer is a=5

OpenStudy (agreene):

I'm saying 1814360/5 +8 = 9! which is true.

OpenStudy (agreene):

which is what you wrote.

OpenStudy (ny,ny):

look on the calculator, its not =)

OpenStudy (lilg132):

did you not understand the way i wrote it out? if you did then i would stick with that method otherwise it may just confuse you more looking at other methods

OpenStudy (agreene):

my calculator says I'm right. http://www.wolframalpha.com/input/?i=solve+a%2F5+%2B+8+%3D+9%21

OpenStudy (ny,ny):

no lilg132, i completely understand lol. when i come up with other questions i will post them. Your answer makes complete sense to me.

OpenStudy (agreene):

lilg132 has a correct solution to a question that is similar--but different from what you asked. I have given you the solution to the question you asked.

OpenStudy (ny,ny):

??? im so confused!!! both questions r the same!! and i dont kno y tht calculator says agreene's answer is correct. If u use the calculator in the start menu located on the desktop, it says lilg132's answer

OpenStudy (ny,ny):

can u plz tell me what is the difference between lilg132's question and yours?

OpenStudy (agreene):

probably because lig132 is answering: a/5 +8 = 9 and I am answering a/5 +8 = 9! and 9 isnt the same as 9!

OpenStudy (ny,ny):

what's 9! ? lol tht was an exclamation mark!! i didnt even know an exclamation mark was a math symbol. lol!

OpenStudy (agreene):

9! is 9*8*7*6*5*4*3*2*1 its 362880 or if you prefer this notation: \[9\times8\times7\times6\times5\times4\times3\times2\times1=362880\] n! = n*(n-1)*(n-2)*...*3*2*1

OpenStudy (ny,ny):

cool!

OpenStudy (lilg132):

hes in like 7th grade i think dont think he needs to know that yet let just stick with the basics :p

OpenStudy (ny,ny):

im a girl =) she's*

OpenStudy (lilg132):

oops my bad *shes :p

OpenStudy (ny,ny):

=)

OpenStudy (agreene):

I learned factorials in probably 3rd or 4th grade--that is a basic.

OpenStudy (ny,ny):

you're saying that you learned my type of question in elementary? I wish i had!